OpenStudy (anonymous):

Write 5x = 40 as a logarithmic function

OpenStudy (anonymous):

IS it \(5^x\) ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Take Log both the sides you will get: \[\log(5^x) = \log(40)\]

OpenStudy (anonymous):

Now use the power rule here: \[\log(m^a) = a.\log(m)\] Use this rule here and tell me what you got..

OpenStudy (klimenkov):

Don't forget the base of the logarithm.

OpenStudy (anonymous):

I am assuming it to be 10..